### v1

@@ 3.1 Fully discrete scheme
hold for all $\left(w_{h},\varphi_{h},\mathbf{v}_{h},q_{h}\right)\in\mathbf{\mathcal{X}}_{h}^{r}$, and $n=0,1,2,\cdots,N-1$, where $\phi_{h}^{0}=R_{h}\phi^{0},~{}\mathbf{u}_{h}^{0}=\mathbf{I}_{h}\mathbf{u}^{0}.$ For simplicity of description, we perform the fully discretization to prove that our numerical scheme satisfies energy stabilization. For the convenience of the following theoretical analysis, we set the parameters $\lambda,M,\nu$ to 1.
3.2 Energy stability
For the above scheme, we can establish the theorem of energy stabilization as follows:
@@ 3.2 Energy stability
Let $\phi_{h}^{n}\in S_{h}^{r}$ and $\mathbf{u}_{h}^{n}\in\mathbf{X}_{h}^{r}$ and the initial values $|\phi_{h}^{0}|\leq C_{1}$ and $E(\phi_{h}^{0},\mathbf{u}_{h}^{0})\leq C_{0}$, i.e., the initial energy is bounded. For all time partitions $\tau$ and mesh size $h$ such that $\tau\leq\min\{Ch/\left(2\|\phi_{h}^{n}\|_{L^{4}}^{2}\right)\}$, the decoupled scheme (3.6)-(3.12) is uniquely solvable and satisfies discrete energy law as follows:
@@
$\lambda\left(\|\nabla\phi_{h}^{n+1}\|^{2}\right.$ $\left.-\|\nabla\phi_{h}^{n}\|^{2}+\|\nabla\phi_{h}^{n+1}-\nabla\phi_{h}^{n}\|^{2}\right)$ (3.21)
$+\frac{2\lambda\tau}{\epsilon^{2}}\left(\frac{1}{4}\delta_{\tau}\|\left(\phi_{h}^{n+1}\right)^{2}-1\|^{2}+\frac{\tau}{4}\left(\|\delta_{\tau}(\phi_{h}^{n+1})^{2}\|^{2}+2\|\phi_{h}^{n+1}\delta_{\tau}\phi_{h}^{n+1}\|^{2}+2\|\delta_{\tau}\phi_{h}^{n+1}\|^{2}\right)\right)$
@@
Recalling inverse inequalities and choosing $\tau\leq Ch/\left(2\|\phi_{h}^{n}\|_{L^{4}}^{2}\right)$, we have
$2\tau\bigg{|}\left(\mu_{h}^{n+1}\nabla\phi_{h}^{n+1},\tilde{\mathbf{u}}_{h}^{n+1}-\mathbf{u}_{h}^{n}\right)\bigg{|}\leq$ $~{}2\tau\|\phi_{h}^{n+1}\|_{L^{4}}\|\nabla\mu_{h}^{n+1}\|\|\tilde{\mathbf{u}}_{h}^{n+1}-\mathbf{u}_{h}^{n}\|_{L^{4}}$ (3.22)
$\leq$ $~{}C\tau\|\phi_{h}^{n+1}-\phi_{h}^{n}+\phi_{h}^{n}\|_{L^{4}}^{2}\|\tilde{\mathbf{u}}_{h}^{n+1}-\mathbf{u}_{h}^{n}\|_{L^{4}}^{2}+M\tau\|\nabla\mu_{h}^{n+1}\|^{2}$
$\leq$ $~{}\frac{1}{2}\|\nabla\phi_{h}^{n+1}-\nabla\phi_{h}^{n}\|^{2}+\frac{C_{2}\tau}{h}\|\phi_{h}^{n}\|_{L^{4}}^{2}\|\tilde{\mathbf{u}}_{h}^{n+1}-\mathbf{u}_{h}^{n}\|^{2}+M\tau\|\nabla\mu_{h}^{n+1}\|^{2},$
@@
$\tilde{E}_{h}^{m+1}-\tilde{E}_{h}^{0}\leq-\tau\sum_{n=0}^{m}\left(2\|\nabla\tilde{\mathbf{u}}_{h}^{n+1}\|^{2}+\|\nabla\mu_{h}^{n+1}\|^{2}\right)-\sum_{n=0}^{m}\left(1-\frac{C_{2}\tau}{h}\|\phi_{h}^{n}\|_{L^{4}}^{2}\right)\|\tilde{\mathbf{u}}_{h}^{n+1}-\mathbf{u}_{h}^{n}\|^{2}.$ (3.23)
In order to the energy to satisfy the law of decay when $m=0$, the following conditions need to be satisfied
$\tau\leq\frac{Ch}{\|\phi_{h}^{0}\|_{L^{4}}^{2}}.$ (3.24)
Since $\|\phi_{h}^{0}\|_{L^{4}}\leq C\|\phi_{h}^{0}\|$ is bounded, the $\tau$ of the above equation (3.24) can be taken to a valid value such that
$\tilde{E}_{h}^{1}-\tilde{E}_{h}^{0}\leq 0.$ (3.25)
So we can obtain that $\|\phi_{h}^{1}\|$ is bounded. When $m=1$, we have
$\tau\leq\min\bigg{\{}\frac{Ch}{\|\phi_{h}^{0}\|_{L^{4}}^{2}},\frac{Ch}{\|\phi_{h}^{1}\|_{L^{4}}^{2}}\bigg{\}},$ (3.26)
then we can dervie that $\|\phi_{h}^{2}\|\leq C$. By analogy, we can obatin the following result
$\tau\leq\min_{0\leq n\leq m}\bigg{\{}\frac{Ch}{\|\phi_{h}^{0}\|_{L^{4}}^{2}},~{}\frac{Ch}{\|\phi_{h}^{1}\|_{L^{4}}^{2}},~{}\cdots,~{}\frac{Ch}{\|\phi_{h}^{n}\|_{L^{4}}^{2}}\bigg{\}}.$ (3.27)
Thus, we can always obtain a positive number $C$ such that $\|\phi_{h}^{n}\|_{L4}\leq C$ is bounded, and derive the desired result (3.13). ∎
@@
Let $\left(\phi_{h}^{n+1},\mu_{h}^{n+1},\tilde{\mathbf{u}}_{h}^{n+1},\mathbf{u}_{h}^{n+1},p_{h}^{n+1}\right)\in\mathbf{\mathcal{X}}_{h}^{r}$ be the unique solution of (3.6)-(3.12). Suppose that $E(\phi_{h}^{0},\mathbf{u}_{h}^{0})<C_{0}$, for any $h,\tau>0$ and $\tau\leq Ch/\left(2\|\phi_{h}^{n}\|_{L^{4}}^{2}\right)$, the following boundedness hold that
@@
where $C_{0}$ is a positive constant dependent on the initial value of $\phi^{0}$ and $\mathbf{u}^{0}$.
@@
Using the lemma 2.1 and 3.1, we have
@@
Using the lemma 2.1 and 3.1, we obtain the result of this lemma. ∎
@@
We note that the bound $\tau\leq Ch/\left(2\|\phi_{h}^{n}\|_{L^{4}}^{2}\right)$ in Lemma 3.1 and 3.2 is attainable, since from (4.26) and Lemma 3.2 we obtain $\|\phi_{h}^{n}\|_{L^{4}}\leq C\|\phi_{h}^{n}\|_{H^{1}}\leq C$. More similar details about energy stabilization under this discrete scheme can be found in reference [4].
@@ 4 Error analysis
In this section, we establish the error analysis for the scheme (3.6)-(3.12). Firstly, we denote $R_{h}:H^{1}(\Omega)\rightarrow S_{h}^{r}$ as the classic Ritz projection [30],
$\|\psi-R_{h}\psi\|_{H^{-1}}\leq Ch^{r+1}\|\psi\|_{H^{r+1}},$ (4.3)
$\|\delta_{\tau}\left(\psi^{n}-R_{h}\psi^{n}\right)\|_{H^{-1}}\leq Ch^{r+1}\|\delta_{\tau}\psi^{n}\|_{H^{r+1}},$ (4.5)
where $\delta_{\tau}$ is defined in (2.7). Let $I_{h}:L^{2}(\Omega)\rightarrow S_{h}^{r}$ and $\mathbf{I}_{h}:\mathbf{L}^{2}(\Omega)\rightarrow\mathbf{X}_{h}^{r}$ are defined the $L^{2}$ projection operators [32, 33] by
$\left(v-I_{h}v,w_{h}\right)=0,~{}\forall w_{h}\in S_{h}^{r},$ (4.6)
$\left(\mathbf{v}-\mathbf{I}_{h}\mathbf{v},\mathbf{w}_{h}\right)=0,~{}\forall\mathbf{w}_{h}\in\mathbf{X}_{h}^{r}.$ (4.7)
$\left\|\mathbf{v}-\mathbf{I}_{h}\mathbf{v}\right\|+h\left\|\nabla\left(\mathbf{v}-\mathbf{I}_{h}\mathbf{v}\right)\right\|\leq Ch^{r+1}\|\mathbf{v}\|_{H^{r+1}}.$ (4.9)
For $\left(\mathbf{v}_{h},q_{h}\right)\in\mathbf{X}_{h}^{r}\times\mathring{S}_{h}^{r-1}$, the Stokes projection [34] $P_{h}$ or $\mathbf{P}_{h}:\mathbf{H}_{0}^{1}(\Omega)\times L_{0}^{2}(\Omega)\rightarrow\mathbf{X}_{h}^{r}\times\mathring{S}_{h}^{r-1}$ are defined by
$\left(\nabla\left(\mathbf{u}-\mathbf{P}_{h}(\mathbf{u},p)\right),\nabla\mathbf{v}_{h}\right)-\left(p-P_{h}(\mathbf{u},p),\nabla\cdot\mathbf{v}_{h}\right)=0,$ (4.10)
$\left(\nabla\cdot\left(\mathbf{u}-\mathbf{P}_{h}(\mathbf{u},p)\right),q_{h}\right)=0.$ (4.11)
To be the simplicity of notations, we denote $P_{h}p=P_{h}(\mathbf{u},p)$ and $\mathbf{P}_{h}\mathbf{u}=\mathbf{P}_{h}(\mathbf{u},p)$.
Lemma 4.1 ([35, 25]).
Under the assumption of regularity (3.1), it holds that
$\|\mathbf{u}-\mathbf{P}_{h}\mathbf{u}\|+h\left(\|\nabla\left(\mathbf{u}-\mathbf{P}_{h}\mathbf{u}\right)\|+\|p-P_{h}p\|\right)\leq Ch^{r+1}\left(\|\mathbf{u}\|_{H^{r+1}}+\|p\|_{H^{r}}\right),$ (4.12)
and
$\|\delta_{\tau}\left(\mathbf{u}^{n}-\mathbf{P}_{h}\mathbf{u}^{n}\right)\|\leq$ $~{}Ch^{r+1}\left(\|\delta_{\tau}\mathbf{u}^{n}\|_{H^{r+1}}+\|\delta_{\tau}p^{n}\|\right.$ (4.13)
$~{}\left.+\|\delta_{\tau}\phi^{n}\|_{H^{r+1}}\|\mu^{n}\|_{H^{2}}+\|\phi^{n-1}\|_{H^{r+1}}\|\delta_{\tau}\mu^{n}\|_{H^{2}}\right).$
where $C$ is the constant the only depends on $\Omega$.
Lemma 4.2 ([36]).
The Stokes projection $\mathbf{P}_{h}(\mathbf{u},p)$ is $\mathbf{H}^{1}$ stable in the case such that
$\|\mathbf{P}_{h}(\mathbf{u},p)\|_{H^{1}}\leq C\left(\|\mathbf{u}\|_{H^{2}}+\|p\|_{H^{1}}\right).$ (4.14)
By defining the discrete Stokes operator $A_{h}=-\mathbf{I}_{h}\Delta_{h}$ [26], it holds
Lemma 4.3 ([25]).
and
Finally, we note the well-known inverse inequalities [25]
$\|v_{h}\|_{L^{4}}\leq Ch^{-1/2}\|v_{h}\|,\quad\forall v_{h}\in S_{h}^{r},$ (4.22)
Lemma 4.4 ([38, 29]).
Suppose $g\in H^{1}(\Omega)$, and $v\in\mathring{S}_{h}^{r-1}$,
$|(g,v)|\leq C\|\nabla g\|\|v\|_{H^{-1}},$ (4.24)
and for all $g\in L_{0}^{2}(\Omega)$, we have
$\|g\|_{H^{-1}}\leq C\|g\|.$ (4.25)
where $C$ is positive constant independent of $\tau$ and $h$.
where $A$ is the continuous Stokes operator [34].
And then, we denote
@@
$e_{\tilde{\mathbf{u}}}^{n+1}=\mathbf{u}^{n+1}-\tilde{\mathbf{u}}_{h}^{n+1}=\mathbf{u}^{n+1}-\mathbf{P}_{h}\mathbf{u}^{n+1}+\mathbf{P}_{h}\mathbf{u}^{n+1}-\tilde{\mathbf{u}}_{h}^{n+1}=\Phi_{\mathbf{u}h}^{n+1}+\Theta_{\tilde{\mathbf{u}}h}^{n+1},$
$e_{\mathbf{u}}^{n+1}=\mathbf{u}^{n+1}-\mathbf{u}_{h}^{n+1}=\mathbf{u}^{n+1}-\mathbf{P}_{h}\mathbf{u}^{n+1}+\mathbf{P}_{h}\mathbf{u}^{n+1}-\mathbf{u}_{h}^{n+1}=\Phi_{\mathbf{u}h}^{n+1}+\Theta_{\mathbf{u}h}^{n+1},$
@@
$R_{1}^{n+1}=\delta_{\tau}R_{h}\phi^{n+1}-\partial_{t}\phi^{n+1},R_{2}^{n+1}=\delta_{\tau}\mathbf{P}_{h}\mathbf{u}^{n+1}-\partial_{t}\mathbf{u}^{n+1},R_{3}^{n+1}=\nabla\left(P_{h}p^{n+1}-P_{h}p^{n}\right),$ (4.35)
@@
Using the (4.26), (4.28), Lemma 4.1 and 4.2, we can obtain
@@
Using the (4.2), and Lemma 4.1, 3.1, 4.5, and 2.1, we can obtain the result of this Lemma. ∎
@@
According to the Lemma 4.1 and 4.2, we can derive
@@
Finally, using the Lemma 4.1, 3.1, 3.2, 4.5, the inequalities (4.2)-(4.5) and the discrete Grönwall’s lemma 2.1, we can ontain the proof of this lemma. ∎
@@
According to the triangle inequality, the inequality (4.3), the Lemma 4.1, 3.1 and 3.2, this completes the proof of Lemma 4.9. ∎
@@
$\sum_{n=0}^{m}\|e_{\mathbf{u}}^{n+1}-e_{\mathbf{u}}^{n}\|_{H^{-1}}^{2}\leq C\tau\left(\tau^{2}+h^{2r}\right).$ (4.84)
@@
According to (4.6) and (3.11), we can obtain $\mathbf{u}_{h}^{n+1}=\mathbf{I}_{h}\tilde{\mathbf{u}}_{h}^{n+1}$. Using the (4.15) and the above results, we have
@@
Squaring both ends of the (4.93) and multiplying the result by $\tau$, summing over $n~{}(0\leq n\leq m)$, and using the Lemma 3.1, 3.2, 4.5, 4.10 and the Theorem 4.2, we can obtain the result of this lemma as follows

### v2

@@ 2 The mathematical prelimilaries
Next, we denote $R_{h}:H^{1}(\Omega)\rightarrow S_{h}^{r}$ as the classic Ritz projection [30],
$\|\psi-R_{h}\psi\|_{H^{-1}}\leq Ch^{r+2}\|\psi\|_{H^{r+1}},$ (2.10)
$\|\delta_{\tau}\left(\psi^{n}-R_{h}\psi^{n}\right)\|_{H^{-1}}\leq Ch^{r+2}\|\delta_{\tau}\psi^{n}\|_{H^{r+1}},$ (2.12)
where $\delta_{\tau}$ is defined in (2.7). Let $I_{h}:L^{2}(\Omega)\rightarrow S_{h}^{r}$ and $\bm{I}_{h}:\mathbf{L}^{2}(\Omega)\rightarrow\mathbf{X}_{h}^{r}$ are defined the $L^{2}$ projection operators [32, 33] by
$\left(v-I_{h}v,w_{h}\right)=0,~{}\forall w_{h}\in S_{h}^{r},$ (2.13)
$\left(\mathbf{v}-\bm{I}_{h}\mathbf{v},\mathbf{w}_{h}\right)=0,~{}\forall\mathbf{w}_{h}\in\mathbf{X}_{h}^{r}.$ (2.14)
$\left\|\mathbf{v}-\bm{I}_{h}\mathbf{v}\right\|+h\left\|\nabla\left(\mathbf{v}-\bm{I}_{h}\mathbf{v}\right)\right\|\leq Ch^{r+1}\|\mathbf{v}\|_{H^{r+1}}.$ (2.16)
For $\left(\mathbf{v}_{h},q_{h}\right)\in\mathbf{X}_{h}^{r}\times\mathring{S}_{h}^{r-1}$, the Stokes projection [34] $P_{h}$ or $\bm{P}_{h}:\mathbf{H}_{0}^{1}(\Omega)\times L_{0}^{2}(\Omega)\rightarrow\mathbf{X}_{h}^{r}\times\mathring{S}_{h}^{r-1}$ are defined by
$\left(\nabla\left(\mathbf{u}-\bm{P}_{h}(\mathbf{u},p)\right),\nabla\mathbf{v}_{h}\right)-\left(p-P_{h}(\mathbf{u},p),\nabla\cdot\mathbf{v}_{h}\right)=0,$ (2.17)
$\left(\nabla\cdot\left(\mathbf{u}-\bm{P}_{h}(\mathbf{u},p)\right),q_{h}\right)=0.$ (2.18)
To be the simplicity of notations, we denote $P_{h}p=P_{h}(\mathbf{u},p)$ and $\bm{P}_{h}\mathbf{u}=\bm{P}_{h}(\mathbf{u},p)$.
Lemma 2.1 ([35, 26]).
Under the assumption of regularity (3.1), it holds that
$\|\mathbf{u}-\bm{P}_{h}\mathbf{u}\|+h\left(\|\nabla\left(\mathbf{u}-\bm{P}_{h}\mathbf{u}\right)\|+\|p-P_{h}p\|\right)\leq Ch^{r+1}\left(\|\mathbf{u}\|_{H^{r+1}}+\|p\|_{H^{r}}\right),$ (2.19)
and
$\|\delta_{\tau}\left(\mathbf{u}^{n}-\bm{P}_{h}\mathbf{u}^{n}\right)\|\leq$ $~{}Ch^{r+1}\left(\|\delta_{\tau}\mathbf{u}^{n}\|_{H^{r+1}}+\|\delta_{\tau}p^{n}\|_{H^{r}}\right).$ (2.20)
where $C$ is the constant the only depends on $\Omega$.
Lemma 2.2 ([36]).
The Stokes projection $\bm{P}_{h}(\mathbf{u},p)$ is $\mathbf{H}^{1}$ stable in the case such that
$\|\bm{P}_{h}(\mathbf{u},p)\|_{H^{1}}\leq C\left(\|\mathbf{u}\|_{H^{2}}+\|p\|_{H^{1}}\right).$ (2.21)
By defining the discrete Stokes operator $A_{h}=-\bm{I}_{h}\Delta_{h}$ [27], it holds
Lemma 2.3 ([26]).
and
Finally, we note the well-known inverse inequalities [26]
$\|v_{h}\|_{L^{4}}\leq Ch^{-1/2}\|v_{h}\|,\quad\forall v_{h}\in S_{h}^{r},$ (2.29)
Lemma 2.4 ([38, 39]).
Suppose $g\in H^{1}(\Omega)$, and $v\in\mathring{S}_{h}^{r-1}$,
$|(g,v)|\leq C\|\nabla g\|\|v\|_{H^{-1}},$ (2.31)
and for all $g\in L_{0}^{2}(\Omega)$, we have
$\|g\|_{H^{-1}}\leq C\|g\|.$ (2.32)
where $C$ is positive constant independent of $\tau$ and $h$.
where $A$ is the continuous Stokes operator [34].
@@ 3.1 Fully discrete scheme
hold for all $\left(w_{h},\varphi_{h},\mathbf{v}_{h},q_{h}\right)\in\mathbf{\mathcal{X}}_{h}^{r}$, and $n=0,1,2,\cdots,N-1$, where $\phi_{h}^{0}=R_{h}\phi^{0},~{}\mathbf{u}_{h}^{0}=\bm{I}_{h}\mathbf{u}^{0}.$ For simplicity of description, we perform the fully discretization to prove that our numerical scheme satisfies energy stabilization. For the convenience of the following theoretical analysis, we set the parameters $\lambda,M,\nu$ to 1.
3.2 Unique solvability and energy stability
For the fully discrete scheme (3.6)-(3.12), we assume that the initial solutions satisfly the following boundedness:
$\left(F(\phi_{h}^{0}),1\right)+\|\phi_{h}^{0}\|_{H^{1}}+\|\mathbf{u}_{h}^{0}\|\leq C,$ (3.13)
where $C$ is a constant independent of mesh size $h$.
In this subsection, we establish the existence of solutions for the fully discrete scheme (3.6)-(3.12). This is achieved by applying the Leray–Schauder fixed point theorem, and we also prove the uniqueness of the numerical solutions. Given that the scheme (3.6)-(3.12) is implemented in a stepwise, decoupled manner, the existence of a solution is demonstrated incrementally. For the Step 2 and Step 3 of the discrete scheme, the proofs of existence and uniqueness are standard results; thus, we omit the detailed discussions here. Instead, our focus is on proving the existence and uniqueness of the solution for Step 1, specifically for the fully discrete scheme (3.6)-(3.7) corresponding to the CH equation.
Lemma 3.1 ([42]).
Let $\mathcal{G}$ denote a compact mapping from a Banach space $B$ into itself, and assume that there exists a constant $C$ such that $\|x\|<C$ for all $x\in B$ and $\alpha\in[0,1]$ satisfying $x=\alpha\mathcal{G}x$. Thus, $\mathcal{G}$ has a fixed point.
Lemma 3.2.
Suppose the initial solutions boundedness (3.13) holds. For given $\tau,h>0$, there exists a solution $\left(\phi_{h}^{n+1},\mu_{h}^{n+1},\mathbf{u}_{h}^{n+1},p_{h}^{n+1}\right)$ to the fully discrete scheme (3.6)-(3.12).
Lemma 3.3.
For $\tau,h>0$ and $\tau\leq C\alpha/{\|\mathbf{u}_{h}^{n}\|_{L^{\infty}}^{2}}$, the fully discrete scheme (3.6)-(3.12) is unique solvability.
The proofs of Lemmas 3.2 and 3.3 are given in Appendix 6.1 and 6.2.
Remark 3.1.
According to the assumption $\tau\leq\frac{\alpha}{C\|\mathbf{u}_{h}^{n}\|_{L^{\infty}}^{2}}$, the above Lemma (3.3) holds. Therefore, we can derive the boundedness of the nth-time-layer velocity as follows:
$\|\mathbf{u}_{h}^{n}\|_{L^{\infty}}\leq C\|\mathbf{u}\|_{L^{\infty}(0,T;\mathbf{L}^{\infty}(\Omega))}\leq C,$ (3.14)
where $C$ is a positive constant independent of $\alpha$, $\tau$, $h$, and $n$. Later in this paper, we will analyze the reasonableness of this condition $\tau\leq\frac{\alpha}{C\|\mathbf{u}_{h}^{n}\|_{L^{\infty}}^{2}}$.
We explain later in Remark 4.1 the feasibility of showing that the inequality holds. For the above scheme, we can establish the theorem of energy stabilization as follows:
@@ 3.2 Unique solvability and energy stability
Let $\phi_{h}^{n}\in S_{h}^{r}$ and $\mathbf{u}_{h}^{n}\in\mathbf{X}_{h}^{r}$ and the initial values $|\phi_{h}^{0}|\leq C_{1}$ and $E(\phi_{h}^{0},\mathbf{u}_{h}^{0})\leq C_{0}$, i.e., the initial energy is bounded. For all time partitions $\tau$ and mesh size $h$ such that $\tau\leq C\alpha/{\|\mathbf{u}_{h}^{n}\|_{L^{\infty}}^{2}}$ and the second step satisfies $\tau\leq Ch/\|\phi_{h}^{n+1}\|_{L^{4}}^{2}$, the decoupled discrete scheme (3.6)-(3.12) is uniquely solvable and satisfies discrete energy law as follows:
@@
$\lambda$ $\left(\|\nabla\phi_{h}^{n+1}\|^{2}-\|\nabla\phi_{h}^{n}\|^{2}+\|\nabla\phi_{h}^{n+1}-\nabla\phi_{h}^{n}\|^{2}\right)$ (3.23)
$+\frac{2\lambda\tau}{\epsilon^{2}}\left(\frac{1}{4}\delta_{\tau}\|\left(\phi_{h}^{n+1}\right)^{2}-1\|^{2}+\frac{\tau}{4}\left(\|\delta_{\tau}(\phi_{h}^{n+1})^{2}\|^{2}\right.\right.$
$\left.\left.\qquad+2\|\phi_{h}^{n+1}\delta_{\tau}\phi_{h}^{n+1}\|^{2}+2\|\delta_{\tau}\phi_{h}^{n+1}\|^{2}\right)\right)$
@@
According to (3.3), we obtain $\|\phi_{h}^{n+1}\|_{H^{1}}\leq C$. Recalling inverse inequality and choosing $\tau\leq Ch$, we have
$2\tau\bigg{|}\left(\mu_{h}^{n+1}\nabla\phi_{h}^{n+1},\tilde{\mathbf{u}}_{h}^{n+1}-\mathbf{u}_{h}^{n}\right)\bigg{|}$ (3.24)
$\leq$ $~{}C\tau\|\nabla\mu_{h}^{n+1}\|\|\phi_{h}^{n+1}\|_{L^{4}}\|\tilde{\mathbf{u}}_{h}^{n+1}-\mathbf{u}_{h}^{n}\|_{L^{4}}$
$\leq$ $~{}\frac{C_{2}\tau}{h}\|\phi_{h}^{n+1}\|_{L^{4}}^{2}\|\tilde{\mathbf{u}}_{h}^{n+1}-\mathbf{u}_{h}^{n}\|^{2}+\tau\|\nabla\mu_{h}^{n+1}\|^{2}$
$\leq$ $~{}\frac{1}{2}\|\tilde{\mathbf{u}}_{h}^{n+1}-\mathbf{u}_{h}^{n}\|^{2}+M\tau\|\nabla\mu_{h}^{n+1}\|^{2},$
@@
$\tilde{E}_{h}^{m+1}-\tilde{E}_{h}^{0}\leq-\tau\sum_{n=0}^{m}\left(2\|\nabla\tilde{\mathbf{u}}_{h}^{n+1}\|^{2}+\|\nabla\mu_{h}^{n+1}\|^{2}\right)-\sum_{n=0}^{m}\frac{1}{2}\|\tilde{\mathbf{u}}_{h}^{n+1}-\mathbf{u}_{h}^{n}\|^{2}.$ (3.25)
Thus, we can always obtain a positive number $C$ that does not mesh size $h$ and time step $\tau$ such that $\|\phi_{h}^{n+1}\|_{L^{4}}\leq\|\phi_{h}^{n+1}\|_{H^{1}}\leq C$ is bounded, and derive the desired result (3.15). ∎
@@
Let $\left(\phi_{h}^{n+1},\mu_{h}^{n+1},\tilde{\mathbf{u}}_{h}^{n+1},\mathbf{u}_{h}^{n+1},p_{h}^{n+1}\right)\in\mathbf{\mathcal{X}}_{h}^{r}$ be the unique solution of (3.6)-(3.12). Suppose that $E(\phi_{h}^{0},\mathbf{u}_{h}^{0})<C_{0}$, for any $h,\tau>0$ and the time step $\tau$ and mesh size $h$ satisfy the conditions of Theorem 3.1, the following boundednesses hold
@@
where $C_{0}$ is a positive constant dependent on the initial values of $\phi^{0}$ and $\mathbf{u}^{0}$.
@@
Using the lemma 2.5 and 3.4, we have
@@
Using the lemma 2.5 and 3.4, we obtain the result of this lemma. ∎
@@
We note that the bound $\tau\leq Ch/\left(2\|\phi_{h}^{n+1}\|_{L^{4}}^{2}\right)$ in Lemma 3.4 and 3.5 is attainable, since from (2.33) and Lemma 3.5 and (6.3), we obtain $\|\phi_{h}^{n+1}\|_{L^{4}}\leq C\|\phi_{h}^{n+1}\|_{H^{1}}\leq C$. More similar details about energy stabilization under this discrete scheme can be found in reference [4].
@@ 4 Error analysis
In this section, we establish the error analysis for the scheme (3.6)-(3.12). We denote
@@
$e_{\tilde{\mathbf{u}}}^{n+1}=\mathbf{u}^{n+1}-\tilde{\mathbf{u}}_{h}^{n+1}=\mathbf{u}^{n+1}-\bm{P}_{h}\mathbf{u}^{n+1}+\bm{P}_{h}\mathbf{u}^{n+1}-\tilde{\mathbf{u}}_{h}^{n+1}=\Phi_{\mathbf{u}h}^{n+1}+\Theta_{\tilde{\mathbf{u}}h}^{n+1},$
$e_{\mathbf{u}}^{n+1}=\mathbf{u}^{n+1}-\mathbf{u}_{h}^{n+1}=\mathbf{u}^{n+1}-\bm{P}_{h}\mathbf{u}^{n+1}+\bm{P}_{h}\mathbf{u}^{n+1}-\mathbf{u}_{h}^{n+1}=\Phi_{\mathbf{u}h}^{n+1}+\Theta_{\mathbf{u}h}^{n+1},$
@@
$R_{1}^{n+1}=\delta_{\tau}R_{h}\phi^{n+1}-\partial_{t}\phi^{n+1},R_{2}^{n+1}=\delta_{\tau}\bm{P}_{h}\mathbf{u}^{n+1}-\partial_{t}\mathbf{u}^{n+1},R_{3}^{n+1}=\nabla\left(P_{h}p^{n+1}-P_{h}p^{n}\right),$ (4.6)
@@
Using the (2.33), (2.35), Lemma 2.1 and 2.2, we can obtain
@@
Using the (2.9), and Lemma 2.1, 3.4, 4.1, and 2.5, we can obtain the result of this Lemma. ∎
@@
According to the Lemma 2.1 and 2.2, we can derive
@@
Finally, using the Lemma 2.1, 3.4, 3.5, 4.1, the inequalities (2.9)-(2.12) and the discrete Grönwall’s lemma 2.5, we can ontain the proof of this lemma. ∎
@@
According to the triangle inequality, the inequality (2.10), the Lemma 2.1, 3.4 and 3.5, this completes the proof of Lemma 4.5. ∎
Remark 4.1.
Finally, we aim to confirm that, as a consequence of $\tau\leq\frac{\alpha}{C\|\mathbf{u}_{h}^{n}\|_{L^{\infty}}^{2}}$, the $a~{}priori$ assumption (6.7) remains valid for all sufficiently small values of $h$. To achieve this, we will show that, for sufficiently small $h$, the following requirement is satisfied. We proceed to prove (6.7) using mathematical induction. Specifically, for $h\leq h_{2}$, (6.7) holds when $n=0$, given that $\mathbf{u}_{h}^{0}=\bm{P}_{h}\mathbf{u}^{0}$. Assuming that (6.7) holds for $n\leq m$, we will demonstrate that this assumption also holds for $n=m+1$. Indeed, if (6.7) holds for $n\leq m$, then (6.6) holds for $n=m$. By using the inverse inequality, for $h\leq h_{2}$, we have
$\left\|\mathbf{u}_{h}^{m+1}\right\|_{L^{\infty}}\leq$ $~{}\left\|\mathbf{u}_{h}^{m+1}-\bm{P}_{h}\mathbf{u}^{m+1}\right\|_{L^{\infty}}+\left\|\bm{P}_{h}\mathbf{u}^{m+1}\right\|_{L^{\infty}}$ (4.55)
$\leq$ $~{}C_{1}h^{-d/2}\left\|\mathbf{u}_{h}^{m+1}-\bm{P}_{h}\mathbf{u}^{m+1}\right\|+\left\|\bm{P}_{h}\mathbf{u}^{m+1}\right\|_{L^{\infty}}$
$\leq$ $~{}C_{1}h^{-d/2}\left\|\mathbf{u}^{m+1}-\mathbf{u}_{h}^{m+1}\right\|+Ch^{-d/2}\left\|\mathbf{u}^{m+1}-\bm{P}_{h}\mathbf{u}^{m+1}\right\|$
$\quad+C\|\mathbf{u}\|_{L^{\infty}\left(0,T;\mathbf{L}^{\infty}(\Omega)\right)}$
$\leq$ $~{}C\|\mathbf{u}\|_{L^{\infty}\left(0,T;\mathbf{L}^{\infty}(\Omega)\right)}.$
Therefore, (6.7) also holds at $n=m+1$. The boundedness explanation for $\|\mathbf{u}_{h}^{n}\|_{L^{\infty}}$ is complete.
@@
$\sum_{n=0}^{m}\|e_{\mathbf{u}}^{n+1}-e_{\mathbf{u}}^{n}\|_{H^{-1}}^{2}\leq C\left(\tau^{2}+h^{2r}\right).$ (4.56)
@@
According to (2.13) and (3.11), we can obtain $\mathbf{u}_{h}^{n+1}=\bm{I}_{h}\tilde{\mathbf{u}}_{h}^{n+1}$. Using the (2.22) and the above results, we have
@@
Squaring both ends of the (4.65) and multiplying the result by $\tau$, summing over $n~{}(0\leq n\leq m)$, and using the Lemma 3.4, 3.5, 4.1, 4.6 and the Theorem 4.2, we can obtain the result of this lemma as follows